# Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) has demonstrated its ability to enhance Large Language Models (LLMs) by integrating external knowledge sources. However, multi-hop questions, which require the identification of multiple knowledge targets to form a synthesized answer, raise new challenges for RAG systems. Under the multi-hop settings, existing methods often struggle to fully understand the questions with complex semantic structures and are susceptible to irrelevant noise during the retrieval of multiple information targets. To address these limitations, we propose a novel graph representation learning framework for multi-hop question retrieval. We first introduce a Multi-information Level Knowledge Graph (Multi-L KG) to model various information levels for a more comprehensive understanding of multi-hop questions. Based on this, we design a Query-Specific Graph Neural Network (QSGNN) for representation learning on the Multi-L KG. QSGNN employs intra/inter-level message passing mechanisms, and in each message passing the information aggregation is guided by the query, which not only facilitates multi-granular information aggregation but also significantly reduces the impact of noise. To enhance its ability to learn robust representations, we further propose two synthesized data generation strategies for pre-training the QSGNN. Extensive experimental results demonstrate the effectiveness of our framework in multi-hop scenarios, especially in high-hop questions the improvement can reach 33.8%. The code is available at: https://github.com/Jerry2398/QSGNN.
1 Introduction
@@
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing the capabilities of Large Language Models (LLMs)  (Peng et al., 2024; Procko & Ochoa, 2024; Gao et al., 2023). By retrieving pertinent information from external knowledge sources and integrating it into the generation process, RAG enables LLMs to ground their responses in factual information, significantly reducing the propensity for hallucinations  (Fan et al., 2024).
A challenging scenario in RAG is handling multi-hop questions, where the answer cannot be directly supported by a single document [^1] Following prior works  (Edge et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), we view document as the retrieval target in this paper. but requires information synthesized from multiple interrelated documents, each contributing a unique piece of the puzzle for the answer  (Dua et al., 2019; Chen et al., 2019). These questions demand that the RAG system accurately identify multiple relevant documents and synthesize them to form a coherent answer. Traditional RAG approaches, which primarily rely on semantic similarity for retrieval  (Karpukhin et al., 2020; Chen et al., 2024; Jiang et al., 2023; Trivedi et al., 2022b), often struggle with multi-hop questions since they overlook the relational dependencies between documents. To overcome this limitation, recent efforts have proposed to leverage Knowledge Graphs (KGs). These methods first construct KGs to model the relationships between documents. During retrieval stage, they either navigate the KGs using graph search strategies  (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2025) or utilize Graph Neural Networks (GNNs)  (Fang et al., 2019; Mavromatis & Karypis, 2024; Luo et al., 2025) to identify relevant information. By using KGs, these approaches can better capture the dependencies between documents, thereby improving the retrieval process for multi-hop questions.
@@
Figure 1: Performance on multi-hop questions. Existing methods struggle to perform well as hop number increases while QSGNN can achieve better performance on high-hop questions.
@@
Although KG based methods have shown improved performance over traditional approaches, they still face significant challenges in addressing the complexities of multi-hop questions. These challenges can be summarized as follows: i) Semantic Comprehensiveness. Multi-hop questions are inherently more complex than one-hop questions from the semantic perspective, requiring a comprehensive understanding of diverse information components. For instance, consider the question: “What was Professor Joe’s opinion about the Apple event in 2023 that featured the announcement of a foldable smartphone concept?” In this case, the entity “Apple” has multiple possible meanings (e.g., a company or a fruit), and the phrase “announcement of a foldable smartphone concept” may refer to a sentence or even a paragraph whose meaning cannot be fully captured by discrete entities within the KG. Such questions not only require the retrieval process to understand context-dependent entities but also demand a full understanding of multi-granular information. However, existing methods primarily focus on entities while ignoring the contextual and multi-granular information. This limitation may lead to the inclusion of irrelevant information (e.g., descriptions of apples as fruits) or the loss of critical details (e.g., the absence of precise entities like “foldable smartphone concept” in the KG), ultimately compromising the performance. ii) Noise Sensitivity. The retrieval process for multi-hop questions is highly sensitive to noise since it has multiple retrieval targets. The noise in any one retrieval target will disrupt the final result. However, existing methods are mostly prone to noise influence. For instance, graph search based methods rely on heuristic strategies to explore relevant information, which is likely to incorporate irrelevant nodes or miss important information. GNN based methods use message passing, which may also aggregate information from irrelevant neighbors. Furthermore, the large search space of KG exacerbates this issue. As shown in Figure  1, the performance of existing methods degrades significantly for high-hop questions, because the number of irrelevant nodes in the KG increases exponentially as the retrieval step increases.
@@
To address these limitations, we propose an effective graph representation learning framework to enhance the RAG performance on multi-hop questions. Our approach is built on three key innovations: i) To comprehensively capture multiple information granularities and their complex relationships, we propose a Multi-information Level Knowledge Graph (Multi-L KG). It consists of three node levels, each representing distinct information granularity. The entity-level captures the basic semantic relationships, the chunk-level models the local contextual relationships, and the document-level represents the global thematic relationships. This design encompasses the fine-grained information relationships and facilitates a more comprehensive understanding for multi-hop questions. ii) To learn representations from the Multi-L KG for multi-hop retrieval, we design a novel GNN model called Query-Specific Graph Neural Network (QSGNN). Specifically, it adopts two kinds of message passing to handle the multiple information components of Multi-L KG: intra-level and inter-level message passing. Intra-level message passing focuses on the basic semantic relationships and the logical coherence relationships within each level, while inter-level message passing considers local-global relationships across levels. In each type of message passing, it considers both semantic relationships and alignment with the query. This design integrates multi-level information and significantly reduces the influence of irrelevant nodes, resulting in high-quality representations for multi-hop retrieval. iii) To further enhance the capability of QSGNN, we propose two data generation strategies to produce synthesized QA pairs for pre-training. This process requires no extra human cost and can be conducted after the construction of Multi-L KG. After pre-training, QSGNN can be quickly adapted to downstream tasks. We conduct extensive experiments to evaluate our method, which demonstrate that QSGNN achieves state-of-the-art performance compared to existing methods, especially in high hop questions the improvement can reach 33.8%.
2 Related Work
Retrieval-augmented generation(RAG). RAG aims at enhancing large language models (LLMs) by retrieving relevant information from external knowledge sources and integrating it into the generation process to enable fact-grounded responses  (Gao et al., 2023; Fan et al., 2024). Traditional RAG methods  (Chen et al., 2024; Li et al., 2023) encode documents from external knowledge sources into dense embeddings and retrieve relevant information by calculating similarity between query and document embeddings. However, these approaches fail to capture the complex relationships between documents, struggling to perform well on the multi-hop questions. In order to adapt the RAG to multi-hop questions, subsequent works  (Jiang et al., 2023; Trivedi et al., 2022b; Su et al., 2024) have proposed multi-step retrieval strategies, where they decompose the original question into multiple sub-questions via LLMs and retrieve relevant information for each sub-question. Although these methods explicitly model the internal structure of the query, they still neglect the multi-level relationships between documents, resulting in limited performance in multi-hop retrieval tasks.
Knowledge Graph for Retrieval-Augmented Generation. Recently, Knowledge Graphs (KGs) have been proposed to facilitate multi-hop retrieval by explicitly modeling complex semantic relationships within and across documents (Peng et al., 2024; Procko & Ochoa, 2024). Current KG-based Retrieval-Augmented Generation (KG-RAG) can be categorized into two main approaches: Graph search based methods and Graph Neural Network (GNN) based methods. Graph search based methods identify seed nodes and leverage heuristic graph search algorithms (such as breadth-first search, community detection, or PageRank) to traverse the KG and retrieve relevant information for multi-hop questions (Edge et al., 2024; Guo et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025). While these methods are intuitive, they are often susceptible to noise due to their dependence on heuristic strategies, which can introduce irrelevant nodes into the retrieval path. Additionally, the large search space of KGs amplifies the impact of noise, leading to information loss or inaccuracies. In contrast, GNN based methods utilize the message passing mechanism of GNNs to directly output retrieval results from the KG in an end-to-end manner (Luo et al., 2025; Wang et al., 2022) or to first identify seed nodes and then retrieve information using graph search algorithms (Mavromatis & Karypis, 2024; Fang et al., 2019; Yasunaga et al., 2021). These methods excel at modeling complex node dependencies through iterative information propagation; however, they also face challenges related to noise, as nodes may aggregate misleading information from irrelevant neighbors. Furthermore, GNN methods often struggle to capture multi-granular semantic information, limiting their effectiveness in addressing complex multi-hop questions.
3 Approach
This section presents the architecture of our proposed method, whose framework is shown in Figure  2. We first introduce the construction of the Multi-information Level Knowledge Graph (Multi-L KG). Next, we describe the detailed design of the Query-Specific Graph Neural Network (QSGNN). Finally, we present the pre-training strategy designed to enhance the performance of QSGNN in multi-hop question retrieval tasks.
@@ 3 Approach
Figure 2: Framework overview. It first constructs Multi-L KG to model the multi-level relationships within corpora. QSGNN is designed to aggregate information from different levels, all the aggregations are guided by query. After pre-training and fine-tuning, it can generate representations for multi-hop questions.
3.1 Multi-information level KG Construction
@@
We define the Multi-L KG as $G=\left(\mathcal{O},\mathcal{C},\mathcal{D},\mathcal{E}_{oo},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cc},\mathcal{E}_{cd}\right)$. It contains three types of node sets: entity set $\mathcal{O}$, chunk set $\mathcal{C}$, document set $\mathcal{D}$ and five types of edge sets: entity-entity $\mathcal{E}_{oo}$, entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$, chunk-chunk $\mathcal{E}_{cc}$, chunk-document $\mathcal{E}_{cd}$. The construction of Multi-L KG involves two main steps:
Node Extraction. Given the document corpus $\mathcal{D}$, we utilize $OpenIE(.)$  (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022) to extract chunks from each document and derive triples from these chunks [^2] The prompt for the OpenIE extraction is detailed in the Appendix  A.10. Note that the term “chunk” can have various meanings and in this paper we view sentence as chunk because we think in most cases sentence is enough to act as a meaningful reference for questions. The subjects and the objects within the triples are collected as entities. The extraction process can be formalized as:
$OpenIE(\mathcal{D})\rightarrow\{(d_{i},\mathcal{C}_{d_{i}},\mathcal{T}_{d_{i}})|d_{i}\in\mathcal{D}\}\rightarrow\mathcal{C}=\{\mathcal{C}_{d_{i}}|d_{i}\in\mathcal{D}\},\mathcal{O}=\{o_{i}|o_{i}\in\mathcal{T}_{d_{i}},d_{i}\in\mathcal{D}\},$ (1)
where $d_{i}$ represents the document $i$, $\mathcal{C}_{d_{i}}$ denotes the set of chunks extracted from $d_{i}$, $\mathcal{T}_{d_{i}}$ is the set of triples derived from $d_{i}$, $o_{i}$ refers to the subject or object within these triples $\mathcal{T}_{d_{i}}$.
Relation Construction. The entity-entity set $\mathcal{E}_{oo}$ is established based on all extracted triples $\{\mathcal{T}_{d_{i}}|d_{i}\in\mathcal{D}\}$, capturing fundamental semantic connections. The chunk-chunk set $\mathcal{E}_{cc}$ is constructed by linking adjacent chunks within a document, as their adjacency reflects a kind of logical coherent expression within a document. The last three edge sets entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$ and chunk-document $\mathcal{E}_{cd}$ are constructed based on containment relationships.
@@
3.2 Query-Specific Graph Neural Network
After constructing the Multi-L KG, the next key challenge is how to effectively retrieve relevant information for multi-hop questions. Existing methods face limitations in their capacity to integrate diverse information levels, and are likely to introduce irrelevant noise due to the complexity of Multi-L KG. To address these challenges, we propose a novel Query-Specific Graph Neural Network (QSGNN). This approach adaptively aggregates information based on the query, effectively mitigating noise while incorporating insights from all information levels within the Multi-L KG, to generate comprehensive, query-specific representations for retrieval. QSGNN has two kinds of message passing process: intra-level message passing and inter-level message passing. These two processes are described in detail as follows:

### v2

# Question-Adaptive Graph Learning for Multi-hop Retrieval Augmented Generation

@@ Abstract.
Retrieval-augmented generation (RAG) has demonstrated its ability to enhance Large Language Models (LLMs) by integrating external knowledge sources. However, multi-hop questions, which require the identification of multiple knowledge targets to form a synthesized answer, raise new challenges for RAG systems. Under the multi-hop settings, existing methods often struggle to fully understand the questions with complex semantic structures and are susceptible to irrelevant noise during the retrieval of multiple information targets. To address these limitations, we propose a novel graph representation learning framework for multi-hop question retrieval. We first introduce a Multi-information Level Knowledge Graph (Multi-L KG) to model various information levels for a more comprehensive understanding of multi-hop questions. Based on this, we design a Question-Adaptive Graph Neural Network (Quest-GNN) for representation learning on the Multi-L KG. Quest-GNN employs intra/inter-level message passing mechanisms, and in each message passing the information aggregation is guided by the question, which not only facilitates multi-granular information aggregation but also significantly reduces the impact of noise. To enhance its ability to learn robust representations, we further propose two synthesized data generation strategies for pre-training the Quest-GNN. Extensive experimental results demonstrate the effectiveness of our framework in multi-hop scenarios, especially in high-hop questions the improvement can reach 33.8%. The code is available at: https://github.com/Jerry2398/QSGNN.
Retrieval-augmented generation; Graph neural networks; Multi-hop Question
journalyear: 2026copyright: ccconference: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 20–24, 2026; Melbourne, VIC, Australiabooktitle: Proceedings of the 49th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’26), July 20–24, 2026, Melbourne, VIC, Australiadoi: 10.1145/3805712.3809693isbn: 979-8-4007-2599-9/2026/07ccs: Information systems Information retrieval query processingccs: Information systems Information retrieval
1. Introduction
@@
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing the capabilities of Large Language Models (LLMs)  (Peng et al., 2024; Gao et al., 2023; Liu et al., 2025, 2023; Hao and Zhang, 2023; Hao et al., 2025; Zhang et al., 2023). By retrieving pertinent information from external knowledge sources and integrating it into the generation process, RAG enables LLMs to ground their outputs in factual evidence, significantly reducing the propensity for hallucinations  (Fan et al., 2024; Procko and Ochoa, 2024).
@@
Figure 1. Performance on multi-hop questions. Existing methods struggle to perform well as hop number increases while Quest-GNN can achieve better performance on high-hop questions.
A challenging scenario in RAG is handling multi-hop questions, where the answer cannot be directly supported by a single document [^1] Following prior works  (Edge et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), we view document as the retrieval unit in this paper. but requires information synthesized from multiple interrelated documents, each contributing a unique piece of the puzzle for the answer  (Dua et al., 2019; Chen et al., 2019). These questions demand that the RAG system accurately identify multiple relevant documents and synthesize them to form a coherent answer. Traditional RAG approaches, which tackle multi-hop questions by decomposing them into a sequence of simpler sub-questions for iterative retrieval (Trivedi et al., 2022b; Fang et al., 2025; Wang et al., 2025; Asai et al., 2024; Yu et al., 2024), often achieve suboptimal performance since they overlook the relational dependencies between documents and may suffer from accumulative intermediate inaccuracies (Hu and Lu, 2024; Arslan et al., 2024). To overcome this limitations, recent efforts have proposed to leverage Knowledge Graphs (KGs). These methods first construct KGs to model the relationships between documents. During retrieval stage, they either navigate the KGs using graph search strategies  (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2025; Chen et al., 2025) or utilize Graph Neural Networks (GNNs)  (Fang et al., 2019; Mavromatis and Karypis, 2024; Luo et al., 2025; Yan et al., 2024; Zhang et al., 2024b; Hu et al., 2025) to identify relevant information. By using KGs, these approaches can better capture the dependencies between documents, thereby improving the retrieval process for multi-hop questions.
@@
Although KG based methods have shown improved performance over traditional approaches, they still face significant challenges in addressing the complexities of multi-hop questions. These challenges can be summarized as follows: i) Semantic Comprehensiveness. Multi-hop questions are inherently more complex than one-hop questions from the semantic perspective, requiring a comprehensive understanding of diverse information components. For instance, consider the question: “What was Professor Joe’s opinion about the Apple event in 2023 that featured the announcement of a foldable smartphone concept?” In this case, the entity “Apple” has multiple possible meanings (e.g., a company or a fruit), and the phrase “announcement of a foldable smartphone concept” may refer to a sentence or even a paragraph whose meaning cannot be fully captured by discrete entities within the KG. Such questions not only require the retrieval process to understand context-dependent entities but also demand a full understanding of multi-granular information. However, existing methods primarily focus on entities while ignoring the contextual and multi-granular information. This limitation may lead to the inclusion of irrelevant information (e.g., descriptions of apples as fruits) or the loss of critical details (e.g., the absence of precise entities like “foldable smartphone concept” in the KG), ultimately compromising the performance. ii) Noise Sensitivity. The retrieval process for multi-hop question is highly sensitive to noise since it has multiple retrieval targets. The noise in any one retrieval target will disrupt the final result. Existing methods are mostly prone to noise. For instance, graph search based methods rely on heuristic strategies to explore relevant information, which is likely to incorporate irrelevant nodes or miss important information. GNN based methods use message passing, which may also aggregate information from irrelevant neighbors. Furthermore, the large search space of KG exacerbates this issue. As shown in Figure  1, the performance of existing methods degrades significantly for high-hop questions, because the number of irrelevant nodes in the KG increases exponentially as the retrieval step increases.
@@
To address these limitations, we propose an effective graph representation learning framework to enhance the RAG performance on multi-hop questions. Our approach is built on three key innovations: i) To comprehensively capture multiple information granularities and their complex relationships, we propose a Multi-information Level Knowledge Graph (Multi-L KG). It consists of three node levels, each representing distinct information granularity. The entity-level captures the basic semantic relationships, the chunk-level models the local contextual relationships, and the document-level represents the global thematic relationships. This design encompasses the fine-grained information relationships and facilitates a more comprehensive understanding for multi-hop questions. ii) To learn representations from the Multi-L KG for multi-hop retrieval, we design a novel GNN model called Question-Adaptive Graph Neural Network (Quest-GNN). Specifically, it adopts two kinds of message passing to handle the multiple information components of Multi-L KG: intra-level and inter-level message passing. Intra-level message passing focuses on the basic semantic relationships and the logical coherence relationships within each level, while inter-level message passing considers local-global relationships across levels. In each type of message passing, it considers both semantic relationships and alignment with the question. This design integrates multi-level information and significantly reduces the influence of irrelevant nodes, resulting in high-quality representations for multi-hop retrieval. iii) To further enhance the Quest-GNN, we propose two data generation strategies to produce synthesized QA pairs for pre-training. This process requires no human cost and can be conducted after the construction of Multi-L KG. After pre-training, Quest-GNN can be quickly adapted to downstream tasks. We conduct extensive experiments to evaluate our method, which demonstrate that Quest-GNN achieves state-of-the-art performance compared to existing methods, especially in high hop questions the improvement can reach 33.8%.
@@
Figure 2. Framework overview. It first constructs Multi-L KG to model the multi-level relationships within corpora. Quest-GNN is designed to aggregate information from different levels, all the aggregations are guided by the question. After pre-training and fine-tuning, it can generate representations for multi-hop questions.
2. Related Work
Retrieval Augmented Generation (RAG). RAG aims at enhancing large language models (LLMs) by retrieving relevant information from external knowledge sources and integrating it into the generation process to enable fact-grounded responses  (Gao et al., 2023; Fan et al., 2024; Hao et al., 2023). Text embedding based methods  (Chen et al., 2024; Li et al., 2023; Liu, 2024) encode documents from external knowledge sources into dense embeddings and retrieve relevant information by calculating similarity between question and document embeddings. However, they fail to capture the complex relationships within the questions and the documents and thus struggle to perform well under multi-hop settings. In order to adapt the RAG to multi-hop questions, subsequent works  (Jiang et al., 2023; Trivedi et al., 2022b; Su et al., 2024; Fang et al., 2025; Wang et al., 2025; Asai et al., 2024; Yu et al., 2024) have proposed multi-step retrieval strategies, where they decompose the original question into multiple sub-questions via LLMs and retrieve relevant information for each sub-question iteratively. Although these methods explicitly model the internal structure of the question, they still neglect the multi-level relationships between documents, resulting in limited performance in multi-hop retrievals.
Knowledge Graph for RAG. Recently, Knowledge Graphs (KGs) have been proposed to facilitate multi-hop retrieval by explicitly modeling complex semantic relationships within and across documents (Peng et al., 2024; Procko and Ochoa, 2024; Zhang et al., 2024a). Current KG-based Retrieval-Augmented Generation (KG-RAG) can be categorized into two main approaches: Graph search based methods and Graph Neural Network (GNN) based methods. Graph search based methods identify seed nodes and leverage heuristic graph search algorithms (such as breadth-first search, community detection, or LLM reasoning) to traverse the KG and retrieve relevant information for multi-hop questions (Edge et al., 2024; Guo et al., 2024; Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025; Sun et al., 2023; Ma et al., 2024; Luo et al., 2023). While these methods are intuitive, they are often susceptible to noise due to their dependence on heuristic strategies, which can introduce irrelevant nodes into the retrieval path. Additionally, the large search space of KGs amplifies the impact of noise, leading to information loss or inaccuracies. In contrast, GNN based methods utilize the message passing mechanism of GNNs to directly output retrieval results from the KG in an end-to-end manner (Luo et al., 2025; Wang et al., 2022) or to first identify seed nodes and then retrieve information using graph search algorithms (Mavromatis and Karypis, 2024; Fang et al., 2019; Yasunaga et al., 2021; Li et al., 2025; He et al., 2024). These methods excel at modeling complex node dependencies through iterative information propagation; however, they also face challenges related to noise, as nodes may aggregate misleading information from irrelevant neighbors. Furthermore, GNN methods often struggle to capture multi-granular semantic information, limiting their effectiveness in addressing complex multi-hop questions.
3. Approach
This section presents the architecture of our proposed method, whose framework is shown in Figure  2. We first introduce the construction of the Multi-information Level Knowledge Graph (Multi-L KG). Next, we describe the detailed design of the Question-Adaptive Graph Neural Network (Quest-GNN). Finally, we present the pre-training strategy designed to enhance the performance of Quest-GNN in multi-hop question retrieval tasks.
3.1. Multi-information level KG Construction
@@
We define the Multi-L KG as $G=\left(\mathcal{O},\mathcal{C},\mathcal{D},\mathcal{E}_{oo},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cc},\mathcal{E}_{cd}\right)$. It contains three types of node sets: entity set $\mathcal{O}$, chunk set $\mathcal{C}$, document set $\mathcal{D}$ and five types of edge sets: entity-entity $\mathcal{E}_{oo}$, entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$, chunk-chunk $\mathcal{E}_{cc}$, chunk-document $\mathcal{E}_{cd}$. The construction of Multi-L KG has two main steps:
Node Extraction. Given the document corpus $\mathcal{D}$, we utilize $OpenIE(.)$  (Angeli et al., 2015; Etzioni et al., 2008; Zhou et al., 2022) to extract chunks from each document and derive triples from these chunks [^2] The prompt for the OpenIE extraction is detailed in the Appendix  A.5. Note that the term “chunk” can have various meanings and in this paper we view sentence as chunk because generally sentence is enough to act as a meaningful reference for questions  (Guo et al., 2024). The subjects and the objects within the triples are collected as entities. The extraction process can be formalized as:
(1) $OpenIE(\mathcal{D})\rightarrow\{(d_{i},\mathcal{C}_{d_{i}},\mathcal{T}_{d_{i}})|d_{i}\in\mathcal{D}\}$
$\mathcal{C}=\{\mathcal{C}_{d_{i}}|d_{i}\in\mathcal{D}\},\mathcal{O}=\{o_{i}|o_{i}\in\mathcal{T}_{d_{i}},d_{i}\in\mathcal{D}\},$
where $d_{i}$ represents the document $i$, $\mathcal{C}_{d_{i}}$ denotes the set of chunks extracted from $d_{i}$, $\mathcal{T}_{d_{i}}$ is the set of triples derived from $d_{i}$, $o_{i}$ refers to the subject or object within these triples $\mathcal{T}_{d_{i}}$, $\mathcal{C}$ denotes the chunk set extracted from $\mathcal{D}$ and $\mathcal{O}$ is the entity set.
Relation Construction. There are five kinds of relationships ($\mathcal{E}_{oo}$, $\mathcal{E}_{cc}$, $\mathcal{E}_{oc}$, $\mathcal{E}_{od}$, $\mathcal{E}_{cd}$) need to be built. The entity-entity set $\mathcal{E}_{oo}$ is established based on all extracted triples $\{\mathcal{T}_{d_{i}}|d_{i}\in\mathcal{D}\}$, capturing fundamental semantic connections. The chunk-chunk set $\mathcal{E}_{cc}$ is constructed by linking adjacent chunks within a document, as their adjacency reflects a kind of logical coherent expression within a document. The last three edge sets entity-chunk $\mathcal{E}_{oc}$, entity-document $\mathcal{E}_{od}$ and chunk-document $\mathcal{E}_{cd}$ are constructed based on containment relationships.
@@
3.2. Question-Adaptive Graph Neural Network
After constructing the Multi-L KG, the next key challenge is how to effectively retrieve relevant information for multi-hop questions. Existing methods face limitations in their capacity to integrate diverse information levels, and are likely to introduce irrelevant noise due to the complexity of Multi-L KG. To address these challenges, we propose a novel Question-Adaptive Graph Neural Network (Quest-GNN). This approach adaptively aggregates information based on the question, effectively mitigating noise while incorporating insights from all information levels within the Multi-L KG, to generate comprehensive, question-adaptive representations for retrieval. Quest-GNN has two kinds of message passing process: intra-level message passing and inter-level message passing. These two processes are described in detail as follows: